John has just turned 22 and is graduating from college. He is thinking about purchasing $200,000 in term life insurance for the coming two years. Calculate the fair premium without expense or profit loadings for John. Assume that the probability of John dying in age 22 to 23, and 23 to 24 are 0.00189 and 0.00200. The claim occurs at the end of the year. The annualized interest rate is 6%. Round to two decimal places when calculating your answer
Answer #1

Probability of John dying in age 22 and 23 , =0.00189

Expected Payment =0.00189*200000=$378

Annualized interest rate =6%=0.06

A. Present Value of Payment =378/(1+0.06)=$356.60

Probability of John dying in age 23 and 24 , =0.00200

Expected Payment =0.00200*200000=$400

Annualized interest rate =6%=0.06

B Present Value of Payment =400/((1+0.06)^2)=$356.00

Fair premium without expense or profit loading=A+B=356.60+356=$712.60